The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $82.80 on December 31, 20Y2.

Marshall Inc.

Comparative Retained Earnings Statement

For the Years Ended December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Retained earnings, January 1

$3,692,000.00

$3,262,000.00

3

Net income

610,000.00

540,000.00

4

Total

$4,302,000.00

$3,802,000.00

5

Dividends:

6

On preferred stock

$10,000.00

$10,000.00

7

On common stock

100,000.00

100,000.00

8

Total dividends

$110,000.00

$110,000.00

9

Retained earnings, December 31

$4,192,000.00

$3,692,000.00

Marshall Inc.

Comparative Income Statement

For the Years Ended December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Sales

$10,860,000.00

$10,000,000.00

3

Cost of goods sold

6,000,000.00

5,460,000.00

4

Gross profit

$4,860,000.00

$4,540,000.00

5

Selling expenses

$2,160,000.00

$2,000,000.00

6

Administrative expenses

1,627,500.00

1,500,000.00

7

Total operating expenses

$3,787,500.00

$3,500,000.00

8

Income from operations

$1,072,500.00

$1,040,000.00

9

Other revenue

99,500.00

20,000.00

10

$1,172,000.00

$1,060,000.00

11

Other expense (interest)

132,000.00

120,000.00

12

Income before income tax

$1,040,000.00

$940,000.00

13

Income tax expense

430,000.00

400,000.00

14

Net income

$610,000.00

$540,000.00

Marshall Inc.

Comparative Balance Sheet

December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Assets

3

Current assets:

4

Cash

$1,050,000.00

$950,000.00

5

Marketable securities

301,000.00

420,000.00

6

Accounts receivable (net)

584,000.00

500,000.00

7

Inventories

430,000.00

380,000.00

8

Prepaid expenses

109,000.00

20,000.00

9

Total current assets

$2,474,000.00

$2,270,000.00

10

Long-term investments

800,000.00

800,000.00

11

Property, plant, and equipment (net)

5,750,000.00

5,184,000.00

12

Total assets

$9,024,000.00

$8,254,000.00

13

Liabilities

14

Current liabilities

$882,000.00

$812,000.00

15

Long-term liabilities:

16

Mortgage note payable, 6%,

$200,000.00

$0.00

17

Bonds payable, 4%,

3,000,000.00

3,000,000.00

18

Total long-term liabilities

$3,200,000.00

$3,000,000.00

19

Total liabilities

$4,082,000.00

$3,812,000.00

20

Stockholders’ Equity

21

Preferred 4% stock, $5 par

$250,000.00

$250,000.00

22

Common stock, $5 par

500,000.00

500,000.00

23

Retained earnings

4,192,000.00

3,692,000.00

24

Total stockholders’ equity

$4,942,000.00

$4,442,000.00

25

Total liabilities and stockholders’ equity

$9,024,000.00

$8,254,000.00

Determine the following measures for 20Y2 (round to one decimal place, including percentages, except for per-share amounts): Assume a 365-day year.

1. Working Capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Number of days' sales in receivables
6. Inventory turnover
7. Number of days’ sales in inventory
8. Ratio of fixed assets to long-term liabilities
9. Ratio of liabilities to stockholders’ equity
10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1

Answer

20Y2 20Y1

1. Working Capital

$1,592,000

$1,458,000

2. Current ratio

2.80

2.80

3. Quick ratio

2.19

2.30

4. Accounts receivable turnover

$20.04

5. Number of days' sales in receivables

$18.2

6. Inventory turnover

14.8

7. Number of days' sales in inventory

24.6

8. Ratio of fixed assets to long-term liabilities

$1.80

$1.73

9. Ratio of liabilities to stockholders' equity

0.8

0.9

10. Times interest earned

$8.88

$8.83

11. Asset turnover

1.3

12. Return on total assets

13.0%

12.8%

13. Return on stockholders' equity

13.0%

14. Return on common stockholders' equity

12.8%

15. Earnings per share on common stock

5.0

4.3

16. Price-earnings ratio

$16.56

17. Dividends per share of common stock

1

1

18. Dividend yield

1.21

Calculation

1. Working Capital

Current Assets - Current Liabilities

20Y2 =2,474,000-882,000 = 1,592,000

20Y1 = 2270000-812000 = 1,458,000

2. Current ratio

Current Assets/Current liabilities

20Y2 =2474000/882000 = 2.80

20Y1 = 2270000/812000 = 2.80

3. Quick ratio

Current Assets Less prepaid expenses and inventory/Current liabilities

20Y2 =(2474000-109000-430000)/882000 = 2.19

20Y1 = (2270000-20000-380000)/812000 = 2.30

4. Accounts receivable turnover

Net Credit Sales/Average Account Receivables

20Y2 =10860000/((584000+500000)/2) = 20.04

5. Number of days' sales in receivables

No. of days in a year/Account receivable turnover ratio

20Y2 = 365/20.03 = 18.2

6. Inventory turnover

Cost of goods sold/ Average Inventory

20Y2 '=6000000/((430000+380000)/2) = 14.8

7. Number of days' sales in inventory

No. of days in a year/Inventory Turnover

20Y2 =365/14.814 = 24.6

8. Ratio of fixed assets to long-term liabilities

Fixed Assets/Long term liabilities

20Y2 '=5750000/3200000 = 1.80

20Y1 '=5184000/3000000 = 1.73

9. Ratio of liabilities to stockholders' equity

Total Liabilities/Shareholders' Equity

20Y2  '=4082000/4942000 = 0.8

20Y1 = 3812000/4442000 = 0.9

10. Times interest earned

earnings before interest & taxes/interest expense

20Y2 '=1172000/132000 = 8.88

20Y1 '=1172000/132000 = 8.83

11. Asset turnover

Net Sales/Average Total assets

20Y2 '=10860000/((9024000+8254000)/2) = 1.3

12. Return on total assets

Earning before interest & Taxes/Total Assets

20Y2 '=1172000/9024000 = 13.0%

20Y1 '=1060000/8254000 = 12.8%

13. Return on stockholders' equity

Net Income/Average Common Stockholders' Equity

20Y2 '=610000/((4942000+4442000)/2) =13.0%

14. Return on common stockholders' equity

Net Income - Preferred dividend/Average Common Stockholders' Equity

20Y2 '=(610000-10000)/((4942000+4442000)/2) = 12.8%

15. Earnings per share on common stock

Net Income Less dividends/No of common equity shares

20Y2 = (610000-110000)/(500000/5) = 5.0

20Y1 = (540000-110000)/(500000/5) = 4.3

16. Price-earnings ratio

Market Value/Earning per share

20Y2 =82.8/5 = 16.56

17. Dividends per share of common stock

Dividend on common stock/No of common stock shares

20Y2 = 100000/(500000/5) = 1

20Y1 '=100000/(500000/5) = 1

18. Dividend yield

Annual Dividend/Market price

20Y2=

Coomon stock = 500000/5 = 1

Market price = 82.0

= 1/82.80 *100 = 1.21